# SupplementSupplementary Information References A Fiber-pigtailed Quantum Dot Device Generating Indistinguishable Photons at GHz Clock-rates

### v1

@@ 6 Photon-Indistinguishability
Next, we investigated the photon-indistinguishability of the emission of the fiber-pigtailed device using two-photon interference experiments to access in a Hong-Ou-Mandel (HOM) setup. For this purpose we interfered consecutively emitted X+ photons in the HOM-setup after spectral filtering via the spectrometer (see S.I., section S2 for experimental details). To quantify the degree of indistinguishability, measurements in co- and cross-polarized configuration are compared. The HOM-experiments are conducted during cooldown #3 for two different temporal delays f 2 ns and 12.5 ns for consecutively emitted photons, providing additional insights in possible dephasing mechanisms [44].
@@ 8 Discussion
In summary, we reported on the fully deterministic fabrication of a directly fiber-pigtailed Purcell-enhanced QD device based on a numerically optimized hCBG microcavity coupled to an UHNA3 single-mode fiber and demonstrate its capability for producing single indistinguishable photons at GHz clock-rates. The achieved Purcell factor of results in short radiative emission lifetimes ps and a strong multiphoton suppression reflected in . The fiber-pigtailed device exhibits photon-indistinguishabilities of 55-80% at 80 MHz excitation repetition rate under quasi-resonant excitation and we demonstrate an in-fiber single-photon coupling efficiency . Moreover, the significant Purcell enhancement enables operation of the fiber-pigtailed device at an excitation clock-rate of 1.28 GHz, resulting in antibunching values and photon-indistinguishabilities under quasi-resonant p-shell excitation. The results presented in this work clearly demonstrate that cutting-edge QD devices with excellent quantum-optical performance can be fully-deterministically integrated with optical SM fibers for the development of robust and practical quantum-light sources. Integrating the fiber-pigtailed device in compact and user-friendly cryocoolers in the future, will enable the implementation of high-performance quantum light source in field-applications of quantum information science.
@@
Another route to improve the degree of coherence under quasi-resonant excitation is a further reduction of the fiber-pigtailed QD-hCBG microcavity system’s -time. While the ps observed in this work for the X+ transition was limited already before the fiber-pigtailing procedure by non-ideal QD-positioning in the hCBG center disc, optimal positioning of the quantum emitter enables 25 and 30 ps as recently reported in Ref. [35]. Modified optimized device designs can additionally increase the Purcell enhancement by a factor of 2-3 [31], while further enhancement might be possible considering the influence of the optical fiber on the cavities -factor, if the fiber-to-hCBG distance can be controlled more precisely. In addition, the further reduced -time would also enhance the temperature-resilience of the photon-indistinguishability enabling the generation of indistinguishable single photons at temperatures achievable with ultra-compact mechanical cryocoolers.ing the obtained fiber-coupling efficiency, the observed strain-attributed wavelength shifts and influences on the emission lines limited the countrates achieved with the fast X+-transition in the current device, but the Mcps countrate of the X- doubtlessly demonstrated the potential of fiber-pigtailed hCBG cavities for highest coupling efficiencies. In this context, additional engineering work is required, for example by minimizing the amount of adhesive, or the membrane area in contact. Alternatively, the strain could be passively countered by an additional adhesion layer on the sample backside, which would exert competing strain [41]. Furthermore, the strain could also be actively controlled by fabricating the device bonded to a piezo substrate [55, 56]. Moreover, it is worth mentioning that the theoretical in-fiber efficiency of 85% is currently limited by the UHNA3 fiber, which is not optimized for the wavelength range around 930-950 nm just above the cut-off wavelength. Moving to telecom O-nd C-Band wavelengths will further increase the hCBGs mode overlap to the UHNA3 optical mode-field, therefore boosting the in-fiber efficiency to above 90% [57]. The shorter wavelength designs would benefit from fibers that have even smaller core diameters.

### v2

@@ 6 Photon-Indistinguishability
Next, we investigated the photon-indistinguishability of the emission of the fiber-pigtailed device using two-photon interference experiments to access in a Hong-Ou-Mandel (HOM) setup. For this purpose we interfered consecutively emitted X+ photons in the HOM-setup after spectral filtering via the spectrometer (see S.I., section S2 for experimental details). To quantify the degree of indistinguishability, measurements in co- and cross-polarized configuration are compared. The HOM-experiments are conducted during cooldown #3 for two different temporal delays of 2 ns and 12.5 ns for consecutively emitted photons, providing additional insights in possible dephasing mechanisms [38].
@@ 8 Discussion
In summary, we reported on the fully deterministic fabrication of a directly fiber-pigtailed Purcell-enhanced QD device based on a numerically optimized hCBG microcavity coupled to an UHNA3 single-mode fiber and demonstrate its capability for producing single indistinguishable photons at GHz clock-rates. The achieved Purcell factor of results in short radiative emission lifetimes ps and a strong multiphoton suppression reflected in . The fiber-pigtailed device exhibits photon-indistinguishabilities of 55-80% at 80 MHz excitation repetition rate under quasi-resonant excitation and we demonstrate a single-photon fiber-coupling efficiency . Moreover, the significant Purcell enhancement enables operation of the fiber-pigtailed device at an excitation clock-rate of 1.28 GHz, resulting in antibunching values and photon-indistinguishabilities under quasi-resonant p-shell excitation. The results presented in this work clearly demonstrate that cutting-edge QD devices with excellent quantum-optical performance can be fully-deterministically integrated with optical SM fibers for the development of robust and practical quantum-light sources. Integrating the fiber-pigtailed device in compact and user-friendly cryocoolers in the future, will enable the implementation of high-performance quantum light source in field-applications of quantum information science.
@@
Another route to improve the degree of coherence under quasi-resonant excitation is a further reduction of the fiber-pigtailed QD-hCBG microcavity system’s -time. While the ps observed in this work for the X+ transition was limited already before the fiber-pigtailing procedure by non-ideal QD-positioning in the hCBG center disc, optimal positioning of the quantum emitter enables 25 and 30 ps as recently reported in Ref. [40]. Modified optimized device designs can additionally increase the Purcell enhancement by a factor of 2-3 [39], while further enhancement might be possible considering the influence of the optical fiber on the cavities -factor, if the fiber-to-hCBG distance can be controlled more precisely. In addition, the further reduced -time would also enhance the temperature-resilience of the photon-indistinguishability enabling the generation of indistinguishable single photons at temperatures achievable with ultra-compact mechanical cryocoolers.
Concerning the obtained , the observed strain-attributed wavelength shifts and influences on the emission lines limited the countrates achieved with the fast X+-transition in the current device, but the Mcps countrate of the X- doubtlessly demonstrated the potential of fiber-pigtailed hCBG cavities for highest coupling efficiencies. To give insight on the repeatability of the coupling technique and the influence of strain, we present experimental data of a second fiber-pigtailed device in S.I., section S10. Also this second device allowed for unambiguous coupling of the respective target QD-hCBG, but showed a plastic deformation of the fiber-to-cavity alignment during its first cooldown, limiting the efficiency. However, the plastic deformation indicates strain relaxation, and in fact this sample did not show the unexpected emission intensity shifts in off-resonant excitation. In this context, additional investigations are required, for example by minimizing the amount of adhesive, or the membrane area in contact. Alternatively, the strain could be passively countered by an additional adhesion layer on the sample backside, which would exert competing strain [48]. Furthermore, the strain could also be actively controlled by fabricating the device bonded to a piezo substrate [30, 42]. Moreover, it is worth mentioning that the theoretical of 85% is currently limited by the UHNA3 fiber, which is not optimized for the wavelength range around 930-950 nm just above the cut-off wavelength. Moving to telecom O- and C-Band wavelengths will further increase the hCBGs mode overlap to the UHNA3 optical mode-field, therefore boosting the fiber-coupling efficiency to above 90% [28]. The shorter wavelength designs would benefit from fibers that have even smaller core diameters.
@@
Note added in proof. - During peer-review of our manuscript, related work by Margaria et al. appeared on arXiv reporting about a fiber-pigtailed micropillar-based single photon source [29].
@@ 12 Disclosure
Supplementary Information
S1: FEM Simulations of FC-QD-hCBGs
| [nm] | [nm] | [nm] | [nm] | [nm] | mRings | [nm] | |
| 360 | 20 | 360 | 100 | 170 | 200 | 5 | 900 |
| 3.460 | 1.450 | 0.12 + 6.33 | 1.493 | 1.451 |
The FEM simulations presented in this work were performed using the commercial FEM software JCMSuite (JCMwave GmbH, 2024). The QD emitter is simulated by a TE dipole source situated in the centre of the central hCBG disc in both vertical and lateral direction. The in-fiber efficiency is calculated as the power obtained from the overlap of emitted CBG emission to the UHNA3-mode profile normalized to the total emitted dipole power. Further details on the simulation setup can be found in [39]. The detailed hCBG-cavity design parameters and used fiber-core radius are listed in Table S1, and the used refractive indices of the involved materials can be found in Table S2.
S2: Influence of fiber distance on simulated Purcell enhancement
As shown in Figure 1 in the main manuscript, the Purcell enhancement of a fiber-pigtailed QD-hCBG cavity depends noticeably on the fiber-to-cavity distance . Figure S2 provides further insights on how the presence of the fiber in close proximity to the hCBG cavity influences the cavity parameters. Figure S2(a) shows the -dependent simulated Purcell factor , quality factor and mode volume relative to the respective values of the hCBG cavity without the present fiber. It is apparent, that varies by up to % from the case without present fiber, while stays nearly constant. It can therefore be assumed that the observed dependency of originates almost exclusively from . Note that the displayed simulation results were obtained from eigenmode simulations of the target hCBG mode in the case of and , while results from scattering simulations at fixed wavelength which takes also contributions of other modes into account. This is the reason for the slight offset of and . It is further found that the mode wavelength exhibits small shifts dependent on , as displayed in Fig. S2.
The simulated -dependency can be understood if the emitted field intensity of the hCBG cavity is considered. Fig. S2(c) shows the field component of the hCBG mode, which shows a near-field pattern with nodes and anti-nodes in vertical direction. If the fiber is placed at the position of minimum field intensity, as is for nm, exhibits a maximum, while gets minimal if the fiber facet lies at positions of high field intensity, such as nm. The reason is that an abrupt change in refractive index, such as the transition of air to fiber-core, is detrimental to mode confinement, a concept that was for example used as "gentle-confinement-designs" for high photonic crystal cavities [4, 52].
We note that the vertical modulation of the hCBG mode’s electric field is present without the fiber present. If a fiber is brought in close proximity in this way, the hCBG’s mode confinement is altered by the induced effective refractive index changes for the given field distribution. The fiber adds additional vertical confinement, effectively becoming part of the hCBG cavity. The predicted noticeable influence of the hCBG mode’s factor by the simulations renders this a versatile method to determine the fiber-to-cavity distance for fabricated structures in the experiment.
Note that Fig. S2(b) also predicts an -dependent wavelength shift of the cavity mode of about nm. The shifts between the cooldowns of the pigtailed QD-hCBG device shown in Fig. 2(c) are however larger than expected from the simulations, indicating that the built-up strain might cause a change of the refractive index as mentioned in the main manuscript.
S3: Experimental setup
Figure S3 shows schematics of the experimental setup used for the quantum-optical measurements of the sample before and after the fiber-pigtailing. For the measurements before the pigtailing in the setup configuration in (a), the QD-hCBG cavity sample is placed in a closed-cycle helium cryostat and cooled to 4 K with a low-temperature compatible microscope objective with NA=0.8 for confocal excitation/emission experiments. Excitation light reaching the sample via the excitation path is either from a pulsed 80 MHz laser system with nominal 2 ps temporal pulse-width (picoEmerald, APE GmbH) or from a broadband white-light source (SLS201L/M, Thorlabs GmbH). The emission from the sample exits the cryostat towards a 0.75 m spectrograph (entrance slit width: 100 m) with 1200 g/mm grating, using a half-wave-plate before the monochromator to match the polarization to the grating. For p-shell experiments, the emission from the sample is filtered with a bandpass filter ( nm) before entering the monochromator. To spectrally resolve the signal, the emission is reflected from the grating onto a 1340 pixel Si-CCD camera cooled to 200 K with a Peltier element. For time-resolved measurements, the emission is reflected from the grating trough a separate spectrograph’s exit (exit slit width: 80 m) into a single mode fiber (SMF 780HP). We expect the spectral filtering caused by the grating and SMF in-coupling to be around 100 eV (0.07 nm). The SMF leads to a Hong-Ou-Mandel setup set as follows: An initial free-space polarization control consisting of half-waveplate, quarter-waveplate and linear polarizer ensures equal splitting at a polarizing BS to match both arms of the Mach-Zehnder interferometer (MZI) in intensity. The arms of the MZI are precisely time-matched for the time-difference ( ns and 12.5 ns for 80 MHz excitation rate, and ps for 1.28 GHz excitation rate). The interference then occurs in a 50:50 fiber-beamsplitter, where the input polarizations are fully-controlled by polarization paddles in fiber to set the co- and cross-polarized interference. The output of the 50:50 fiber-beamsplitter leads to superconducting nanowire single photon detectors (SNSPDs) (Single Quantum EOS CS, Single Quantum B.V.) connected to time tagging electronics (QuTag, qtools GmbH). Additional polarization paddles on the fibers leading to the detector match the polarization to the nanowire orientation. The correct setting of polarization and time delay in the MZI is confirmed by interfering laser pulses at the QD emission wavelength and observing close to 100 interference contrast.
For the measurements of the FC-device, the sample was placed in the cryostat and the 780HP fiber output of the UHNA3-780HP fiber patchcord was connected to the cryostats vacuum fiber-feedtrough. On the room temperature side of the feedtrough, a 90:10 fiber-based beamsplitter was connected, whose 10% input arm was used to connect the above mentioned 80 MHz pulsed laser for optical excitation, or the broadband white-light source for reflection measurements. The 90% fiber-beamsplitter output was connected to a free-space out-coupler towards the spectrometer. The same spectrometer and Si-CCD was used as for the free-space measurements, but was entered from an alternative entrance.
The efficiency estimation for the FC setup is discussed in S.I., section S8.
S4: Fiber-pigtailed spectra during cooldown
Figure S4 shows the temperature dependent spectra during the three cooldowns of the FC-device for an excitation wavelength of nm. For cooldown #1, the excitation power conditions at the QD were unknown, and the excitation power and integration time on the CCD was adjusted in the range above 54 K to check for QD state saturation. The changes in intensity above 54 K stem from these excitation power variations, and not from the sample itself. Once the cooldown was started, it could not be stopped until base temperature was reached, causing the intensity fluctuations to be recorded.
Fig. S4(d)-(f) show the temperature-dependent emission of the X- line as a close-up for the respective cooldowns. The dotted line follows equation (1) for a temperature dependent band gap energy of GaAs (shifted by a set confinement energy for each cooldown), with the corresponding parameters listed in [35]:
The good agreement with theory indicates that the FC-device reaches the indicated temperatures (ultimately 4.8 K), and the wavelength shifts compared to the sample before pigtailing cannot be explained by temperature, as elaborated in the main text. We note further that equation (1) with InAs parameters does not fit the experimental data well, indicating that the QD’s temperature dependence is either largely determined by the surrounding GaAs matrix material, or significant Ga incorporation takes place during the growth.
S5: Assigment of QD states to emission lines.
To assign the observed emission lines in the spectrum of the fiber-pigtailed QD-hCBG device excitation power dependent and detection polarization dependent measurements were conducted during during the first cooldown. The data is displayed in Figure S5, with the assigned states for the spectrum shown in Fig. S5(a).
The XX and X emission lines are assigned by the visible fine-structure splitting in the polarization resolved detected emission intensity in Fig. S5(c), and X is distinguished from XX based on the lower slope from the excitation power dependent measurements in Fig. S5(b).
Neither of the assigned X+, XX+ and X- states show a (resolvable) fine-structure splitting for varying polarization, indicating trionic states. X+ and XX+ are distinguished based on their power dependency. The appearance of X+ and XX+ at shorter wavelength (i.e. anti-binding) compared to the X is commonly observed for InAs QDs, while the negative trion appears at longer wavelengths (i.e. binding) compared to X [41].
S6: Off-resonant spectra before and after fiber-pigtailing
Figure S6 shows the spectra in Fig. 2(a) in the main text under off-resonant excitation ( nm) before and after the fiber-pigtailing plotted in logarithmic scale. The wavelength on the x-axis is normalized to the emission wavlength of the X+ transition. Respective QD states corresponding to emission lines are indicated, as well as the filtered-out spectral region by the employed bandpass filter for the spectrum before the coupling.
X+, XX, XX and X lines appear at identical spectral positions, clearly proving that the fiber-pigtailing was deterministic for the pre-selected QD-hCBG cavity. The strong emission background in the spectrum before the pigtailing in Fig. S6(a) is caused by 10 higher excitation power . Similarly, was slightly higher for cooldowns #1 and #4 for the displayed spectra, causing slightly varying backgrounds. An excition wavlength-dependent comparison before and after the pigtailing is discussed further in S.I., section S7.
S7: Photoluminescence excitation before and after fiber-pigtailing
Figure S7 shows photoluminescence excitation (PLE) data before and after the pigtailing, with detailed PLE scans before in Fig. S7(a) and after the pigtailing for cooldown in in Fig. S7(b). The respective p-shell and X+ wavelengths are indicated. Fig. S7(c) shows the intensity of the X+ line for varying excitation wavelengths normalized to the trion emission wavelengths, i.e. the s-p splitting. The fact that identical s-p splitting is observed before and after the pigtailing proves the deterministic fiber-pigtailing technique.
S8: Setup efficiency estimation
Figure S8 shows the measurement to estimate the setup efficiency for the characterisation of the fiber-pigtailed sample. The fiber-based beamsplitter’s input, which would be connected to the cryostat’s fiber-feedthrough for the characterization of the FC-device, is connected to a continuous-wave (cw) laser at with input power W. The 10% output of the fiber-based beamsplitter is blocked, and the cw-laser signal reaches the spectrometer via the 90% output. The power is measured at the entrance of the spectrometer. The power is the cw-laser power at the exit of the spectrometer, and is attenuated with OD-filters before entering the fiber towards the SNSPD detectors. is the observed countrate on the SNSPDs.
We find = 0.327, = 0.333 and = 0.45. This leads to an overall setup efficiency from the fiber-based beamsplitter input to the detectors of 0.049. For the experimental in-fiber efficiency at the UHNA3 fiber-facet, the efficiency by the splice (0.95) and by the vacuum fiber-feedthrough (0.6) have to be additionally taken into account, as described in the main text.
S9: Influence of on
The following S.I. provides a reference for the measured two-photon-indistinguishablity values in p-shell excitation for the fiber-pigtailed device, since we did not measure before the pigtailing. The -influence on the indistinguishability of two emitted photons can be described as [19, 32]:
with the complementary error-function , factor , and the inhomogeneously broadened QD emission line following a Gaussian distribution with FWHM-linewidth . Figure S9 shows the measured values of the pigtailed QD-hCBG device for p-shell excitation at different fractions of discussed in the main text plotted against the observed -time. The results of equation (2) are shown for -values between 1.3 GHz and 6 GHz (i.e., linewidths of 0.005 nm and 0.028 nm, respectively), in agreement with inhomogenous-broadening values previously reported for InAs QD-hCBG cavities under p-shell excitation [40]. The photon-indistinguishability observed for the fiber-pigtailed device, as discussed in the main text, are hence very similar to free-space operated QD-hCBG cavities. It can be expected, that even higher indistinguishabilities under non-resonant excitation can be reached, if can be further reduced.
S10: Additional data from a second fiber-pigtailed QD-hCBG device
We performed a fiber-pigtailing process for a second sample. In the following, we will refer to this second pigtailed as FC-CBG #2. In contrast to the discussed pigtailed device in the main manuscript, device FC-CBG #2 showed a sudden decrease in emission intensity a couple of minutes after reaching base temperature in a first cooldown. The emission properties observed during a second cooldown are displayed in Figure S10.
Figure S10(a) shows recorded spectra under off-resonant excitation during the second cooldown. Figure S10(b) shows the device’s emission at K in green, as well as the spectrum before the pigtailing in blue. Despite the reduced intensity after the first cooldown, the fiber-to-cavity alignment was not completely lost, and further spectra, photoluminescence excitation (PLE) scans and time-resolved measurements could be conducted during the second cooldown, as displayed in Figure S10(c-f). Based on the observed quasi-resonances in the PLE and the spectral positions of the assigned QD states, the pigtailed QD-hCBG device is unambiguously identified as the target device before the pigtailing, albeit again with a slight red shift. The measured -time of X+ appears to be once again slower than before the pigtailing. Note that for this device FC-CBG #2, it was not possible to obtain information on the cavity mode from reflection measurements, since the signal-to-noise ratio was too low stemming from the fiber-displacement, so no information can be gained on the mode’s -factor after the pigtailing.
To put this observed properties of device FC-CBG #2 in perspective with the pigtailed device in the main manuscript: we believe that the pigtailed device discussed in the manuscript experienced an elastic deformation and strain built-up during the first thermal cycle, which appears to equilibrated after the first cooldown. For device FC-CBG #2, the thermal stressed appeared to be enough to cause a plastic deformation of the fiber-to-hCBG connection, which resulted in the observed deteriorated alignment and loss in emission intensity. The plastic deformation however, is also believed to have caused a partial relaxation of the strain accumulated during the first cooldown. This can explain why there are no intensity shifts observed in this second sample, compared to the device discussed in the manuscript.
The reason why device FC-CBG #2 showed a plastic deformation, while the device discussed in the manuscript had intact fiber-alignment after the first cooldown could originate from several factors: Firstly, the temperature ramp of the cryostat during 1st cooldown of device FC-CBG #2 was steeper than for the second cryostat used for the later cooldowns, potentially causing higher thermal stress. Secondly, we think that the membrane area being in contact with the fiber might play a crucial part in the strain distribution and potential deformations.